OpenStudy (anonymous):

The one-to-one function h is defined by h(x)=-8x-7/-5-3x Find h^-1 , the inverse of h . Then, give the domain and range of h^-1 using interval notation.

OpenStudy (anonymous):

f(x) = (9x+8) / ( -9+8x) y = (9x+8) / ( -9+8x) to find the inverse 1. replace xby y and yby x 2. solve for y 3. replace y by f^-1(x) 1. x=( 9y+8) / (-9+8y) 2. -9x +8xy = 9y +8 8xy-9y = 8+9x y(8x-9) = (8+9x) y = (8+9x) / (8x-9) 3. f^-1(x) = (8+9x) / (8x-9) domain : set of all real values except x=9/8) range: set of all real values except 9/8 ( f(x) cannot =9/8) because in this graph there is a vertical asymptotes at x=9/8and horizontal asymptotes at y=9/8

OpenStudy (anonymous):

@pratu043 thanx

OpenStudy (anonymous):

h^-1 (x)=? domain (h^-1)=? range (h^1)=?